# Force Prompting: Video Generation Models Can Learn and Generalize Physics-based Control Signals

### v1

@@ 1 Introduction
Humans develop an intuitive understanding of how objects respond to forces since infancy (Wilkening and Cacchione, 2010; Ullman et al., 2017): a gentle poke causes a plant to sway, while a breeze creates rippling patterns across fabric. Do video generation models, which encode powerful visual and motion priors through internet-scale pretraining, possess a similar level of intuitive physics understanding? And if so, how to elicit their capabilities to interact with force inputs? A positive answer to these questions would provide a more flexible and expressive interface for video content creation, enable interactive video generation with user input (e.g., generating a video game), and eventually lead to an intuitive world model for intelligent agents to plan and make decisions with.
@@
We implement Force Prompting by introducing additional force control as local or global vector fields on a video generation model (Yang et al., 2024) conditioned on initial frame and text. We also curate an evaluation benchmark of diverse objects and motion types to evaluate global and local force prompts. As illustrated in Figure 1, our main finding is that despite the synthetic visual appearance and few objects (flying flags and rolling balls) in our training data, video generation models can indeed learn to execute fine-grained force prompts, and exhibit surprisingly strong generalization behavior across diverse settings, object shapes and materials, geometry, and affordances. Through extensive human evaluations, we demonstrate that Force Prompting exhibits superior adherence to physical instruction while maintaining realistic motion and visual quality, when compared to text-conditioned baselines. This validates our hypothesis that synthetic data can teach video generation models intuitive physics and control without damaging their video priors. We further show that simply extrapolating the future by treating forces as local trajectories is insufficient, and our approach significantly outperforms the state-of-the-art in trajectory-controlled video generation (Geng et al., 2024). Notably, Force Prompting can be trained in approximately a single day on four NVIDIA A100 GPUs. We also try to understand the cause of this strong generalization and perform a careful ablation on the training data. We find two elements that appear important to generalization: visual diversity in the training data with respect to the control signal and the usage of certain text keywords at training time, which appear to help elicit the understanding of force control signals.
@@ 2 Related Works
Video generation: In the last several years, video generation models have made rapid progress in visual quality and realistic dynamics (Singer et al., 2022; Ho et al., 2022; Blattmann et al., 2023; Girdhar et al., 2023; Bar-Tal et al., 2024; Brooks et al., 2024). In particular, Sora (Brooks et al., 2024) was one of the first video generation models to demonstrate truly compelling diverse real-world physical phenomena and directly advocated for the future use of using video generators as simulators for the physical world. In the last half year, significant progress has been made by open source models such as CogVideoX (Yang et al., 2024) and Wan 2.1 (Wang et al., 2025), even approaching the quality of closed-source models. While these models act as strong video priors, they primarily use text and images as input and lack precise control over general actions or other physical inputs.
@@
Controllable video generation: As video models have rapidly progressed, so too has the accompanying field of controllability for these models with the majority of work in this domain focusing on either camera control (He et al., 2024; Zheng et al., 2024; Sun et al., 2024) or various paradigms of motion control (Yin et al., 2023; Chen et al., 2023; Wang et al., 2024; Shi et al., 2024; Niu et al., 2024; Wu et al., 2024; Geng et al., 2024; Li et al., 2024a; Namekata et al., 2024; Zhang et al., 2024b), such as drag-based, trajectory-based, and optical flow-based techniques. Many of the existing motion control models Yin et al. (2023); Chen et al. (2023); Zhang et al. (2024b) require the complete pre-specified trajectory, specifying the location of the pixel on every generated frame. This reliance on full temporal information makes it difficult to use these models for simulation or prediction tasks.
@@
Motion Prompting (Geng et al., 2024), a concurrent work, uses spatio-temporally sparse trajectories as a conditioning signal, enabling users to specify motion over a few frames for video extrapolation. While this might superficially resemble force control, crucial distinctions exist. First, global phenomena like wind or fluid dynamics are naturally expressed as forces but are difficult or impossible to represent with trajectories. Second, applied forces fundamentally depend on an object’s mass or material properties - a dependency absent when specifying motion or location (e.g., identical forces induce greater displacement in lighter objects). Third, specifying an object’s location across a few frames is not equivalent to an applied force; the same observed motion could result from numerous alternative causes, such as camera movement or internal object changes. We compare to Motion Prompting and demonstrate significantly better adherence to the conditioning force.
@@
Interactive world models: Paralleling the interest in video generation models, interactive world models (Ha and Schmidhuber, 2018) have gained significant attention. Despite extensive research in this area, investigations have predominantly concentrated on video game environments Valevski et al. (2024); Che et al. (2024); Bruce et al. (2024). While a few contemporary studies have begun exploring real-world applications (e.g., Bar et al. (2024); Agarwal et al. (2025)), none explores interactions besides camera control or text. In contrast, our work focuses on interaction through physical forces.
@@
Physical simulators and hybrid approaches: Early work (Davis et al., 2015a, b) on generating video based on intuitive forces extracts modal bases of vibrating objects in 2D image space; these works, as well as their modern adaptations (Li et al., 2024b), represented motion as a series of vibrations with different frequencies and intensities, which works well for vibration-like motions but struggles to represent many types of motion, such as linear motion. This led to an alternative research direction explicitly incorporating physics solvers (Chen et al., 2022; Zhong et al., 2024; Le Cleac’h et al., 2023; Xie et al., 2024; Zhang et al., 2024a; Huang et al., 2024; Liu et al., 2024a; Lin et al., 2024; Aira et al., 2024). However, almost all of these techniques require the 3D geometry of the scenes. Recent work has focused on combining both physics simulators and generative models, trying to get the best of both worlds: accurate dynamics from the simulator and better appearance from generative models. For example, PhysGen (Liu et al., 2024b) uses a rigid-body physics solver to model object collisions and then renders these scenes through a video generator, and PhysMotion (Tan et al., 2024) uses a combination of a 3D physics solver and a video generation model. However, due to their usage of physics simulators, they are limited in the types of dynamics they can model. In contrast, we explore using the video generation as a simulator and do not use a physics simulator at inference time. Concurrent work (Li et al., 2025) also explores the use of simulated videos to finetune generative models, but their focus is on modeling object freefall as opposed to learning physics-based control.
@@ 3.1 Synthetic training data
To construct our global wind force dataset, we use a physics simulator to generate videos of flags waving in the wind. And we construct our local point force dataset in two parts: for the first part, we use a physics simulator to generate videos of a ball rolling across the ground; and for the second part, we use a model (Zhang et al., 2024a) which integrates 3D Gaussians and a physics simulator to generate videos of a plant being poked. We provide more detail below.
Global force dataset: We use Blender to construct a dataset of flags waving in response to varying wind conditions. In order to generate a diverse dataset, we randomize multiple parameters for each video: flag quantity ($\mathrm{Unif}\{1,\dots,64\}$), flag color (from a set of 100), flag positions, camera placement, HDRIs, which are 360-degree panoramic images used for lighting and background purposes (selected from 50 options on Polyhaven), wind direction in $[0,360)$, and wind speed in $[0,1]$, where 0 corresponds to no wind, and $1$ corresponds to very strong wind. Each video captures the flags’ transitions from stationary to wind-affected state. Our training dataset has 15k videos.
Local force dataset: The first scenario in our dataset comprises 12k videos of balls, with one of them rolling in response to being pushed by an unseen point-wise force (the force actor is not rendered), and the other balls remaining stationary. We generate these videos using Blender with randomized parameters: ball quantity ($\mathrm{Unif}\{2,3,4\}$), ball textures (soccer balls using a Polyhaven mesh [$p=2/3$] , or bowling balls modeled as smooth spheres [$p=1/3$]), ball colors (from a set of 108), ball positions, camera position, ground textures (from 42 Polyhaven options), target ball selection, force angle in $[0,360)$, and force magnitude in [0,1]. We assign the bowling ball to be four times the mass of the soccer ball with the goal of teaching the model mass-based dynamics. The second scenario (11k videos) utilizes PhysDreamer (Zhang et al., 2024a), a generative-simulator hybrid, and features videos of a carnation swaying back and forth in response to being poked by an unseen force. We generate these videos with randomized camera position, contact points, force angles, and magnitudes. We use a mixed dataset with the goal of teaching the model that a point force can result in both simple linear motion, and complex oscillatory dynamics, depending on what type of object the force is applied to. In both scenarios, the force magnitude 0 corresponds to a very gentle poke, and the force magnitude $1$ corresponds to a much stronger poke.
@@ 3.2 Local and Global Force Prompts
Figure 3: Qualitative results for the Local Force (Poke) model. Top section: For local forces, the control signal can specify both the location, magnitude, and direction of the force. Bottom section: despite the limited training data, the model generalizes to different types of motion.
@@
We note that force values across the ball rolling and plant poking training videos are not calibrated to any absolute physical scale. Instead, they follow intuitive relative physics where smaller force values (approaching $F=0$) correspond to gentle pokes resulting in minimal initial displacement, while larger force values produce stronger pokes with correspondingly greater initial displacement. We also wish to highlight that our force prompting models are fundamentally different from video generative models with trajectory-based control such as (Zhang et al., 2024b; Geng et al., 2024). This is because the gaussian blob which serves as the force indicator for the point force model is generally far away from the pixels that it affects, as demonstrated in the complex oscillatory motion of the swaying flower in Figure 2. Similarly, the wind force control signal under-specifies which points must move to which locations, as that control signal is global and causal.
@@ 3.3 Architecture and Training
We build the force prompting models on top of CogVideoX-5B-I2V (Yang et al., 2024), a video generative model which accepts text and initial frame as conditional inputs. This model generates 49-frame videos at 8-fps. In order to integrate force prompt conditioning, we add a ControlNet (Zhang et al., 2023) which inputs a physics control prompt $\pi$, processing it through downscaling, encoding, and temporal compression before combining with hidden states via a zero convolution. The ControlNet clones the first six transformer layers and fine-tunes them while keeping the base model’s transformer layers frozen. We base our implementation on (Karachev and Xu, 2025) with modifications to adhere more closely to the original ControlNet design. We train the models on a four 80 GB A100 GPU cluster for 5000 training steps, which takes approximately one day. Training uses an instantaneous batch size per device of $1$, with two gradient accumulation steps, for an effective batch size of $8$. Full hyperparameter details are listed in Appendix A.1.
@@ 4 Quantitative and Qualitative Results
Baseline models: The first baseline is text-only, zero-shot, which uses the original CogVideoX model and describes the intended force with a string and appending it to the end of the original text prompt. Two example prompt string suffixes are “the apple is moved very forcefully, upwards and to the left”, and “the wind is medium strength, blowing right”. The second baseline is text-only, fine-tuned, which has the same ControlNet architecture as our force prompting model, but with zero-tensor control signals, as well as force suffixes added to the end of the text prompts during training. Our third baseline is Motion Prompting (Geng et al., 2024), built on Lumiere (Bar-Tal et al., 2024) (run by the authors). It is the only track-conditioned model that accepts temporally sparse tracks as conditioning signal. We simulate force prompt’s impulse by tracing push paths from target objects for the first 3 frames. While the model is meant to accept temporally sparse trajectories, 3 frames of trajectory is out of domain for the intended use case of Motion Prompting.
Human study for local force benchmark: We create a benchmark by curating 63 images from Pexels demonstrating three categories of physical interactions: 1) linear movement patterns (toy car, toy train on straight track, hot air balloon); 2) oscillatory movement patterns (windmill, pendulums, ornament, and swing); and 3) complex movement patterns (toy train on circular track, various plants including ivy, apple tree, and flowers). Table 1 presents human evaluation results for point forces, showing that despite training only on ball rolling (linear) and plant poking (complex) scenarios, our force prompting model demonstrates strong generalization across all motion categories. We visualize some of these generalization patterns in Figure 3.
@@
Human study comparing to PhysDreamer: The point force model, trained on data from a single carnation, demonstrates remarkable generalization to other plants, as we illustrate in Figures 1 and 3. To evaluate this generalization quantitatively, we compare our approach against PhysDreamer (Zhang et al., 2024a), which employs 3D assets and an integrated physics simulator. Using their benchmark dataset of six plant species, our results in Table 2 show that the point force model successfully generalizes to various roses, tulips, and alocasia without specific training on these plants. While we do not claim to replace physics-based simulation approaches, our purely neural method offers exceptional generalizability and produces responses that align with “intuitive physics,” effectively conveying plausible physical interactions to human evaluators. We also also conduct an extended qualitative comparison with 8 other physics simulation models; see Appendix B.2 for more details.
@@ 5.1 Ablation Study #1: Composition of Synthetic Dataset
Global force training dataset ablation: For the global wind force task, we evaluate two diversity factors: flag quantity and background variety. We find that training with a single background leads to models that follow force physics but frequently fail to differentiate between foreground and background, reducing visual quality. Similarly, when restricting scenes to contain only one flag instead of a variable number ($\mathrm{Unif}\{1,\dots,64\}$), the model successfully models cloth mechanics but fails to generalize to other materials. In these cases, smoke from campfires remains unaffected by wind, and confetti either doesn’t respond or stays unnaturally suspended. We also observe that bubbles don’t respond to wind, while human limbs incorrectly billow like cloth. These failures indicate that insufficient scene diversity causes the model to overfit to stationary backgrounds and limited material interactions. These findings are illustrated in Figure 7 as well as the project webpage.

### v2

@@ 1 Introduction
Humans develop an intuitive understanding of how objects respond to forces since infancy (wilkening2010children; ullman2017mind): a gentle poke causes a plant to sway, while a breeze creates rippling patterns across fabric. Do video generation models, which encode powerful visual and motion priors through internet-scale pretraining, possess a similar level of intuitive physics understanding? And if so, how to elicit their capabilities to interact with force inputs? A positive answer to these questions would provide a more flexible and expressive interface for video content creation, enable interactive video generation with user input (e.g., generating a video game), and eventually lead to an intuitive world model for intelligent agents to plan and make decisions with.
@@
We implement Force Prompting by introducing additional force control as local or global vector fields on a video generation model (yang2024cogvideox) conditioned on initial frame and text. We also curate an evaluation benchmark of diverse objects and motion types to evaluate global and local force prompts. As illustrated in Figure 1, our main finding is that despite the synthetic visual appearance and few objects (flying flags and rolling balls) in our training data, video generation models can indeed learn to execute fine-grained force prompts, and exhibit surprisingly strong generalization behavior across diverse settings, object shapes and materials, geometry, and affordances. Through extensive human evaluations, we demonstrate that Force Prompting exhibits superior adherence to physical instruction while maintaining realistic motion and visual quality, when compared to text-conditioned baselines. This validates our hypothesis that synthetic data can teach video generation models intuitive physics and control without damaging their video priors. We further show that simply extrapolating the future by treating forces as local trajectories is insufficient, and our approach significantly outperforms the state-of-the-art in trajectory-controlled video generation (geng2024motion). Notably, Force Prompting can be trained in approximately a single day on four NVIDIA A100 GPUs. We also try to understand the cause of this strong generalization and perform a careful ablation on the training data. We find two elements that appear important to generalization: visual diversity in the training data with respect to the control signal and the usage of certain text keywords at training time, which appear to help elicit the understanding of force control signals.
@@ 2 Related Works
Video generation: In the last several years, video generation models have made rapid progress in visual quality and realistic dynamics (makeavideo2022; imagenvideo2022; svdblattmann2023; emuvideo2023; bar2024lumiere; sora2024). In particular, Sora (sora2024) was one of the first video generation models to demonstrate truly compelling diverse real-world physical phenomena and directly advocated for the future use of using video generators as simulators for the physical world. In the last half year, significant progress has been made by open source models such as CogVideoX (yang2024cogvideox) and Wan 2.1 (wan2025), even approaching the quality of closed-source models. While these models act as strong video priors, they primarily use text and images as input and lack precise control over general actions or other physical inputs.
@@
Controllable video generation: As video models have rapidly progressed, so too has the accompanying field of controllability for these models with the majority of work in this domain focusing on either camera control (he2024cameractrl; zheng2024cami2v; sun2024dimensionx) or various paradigms of motion control (yin2023dragnuwa; chen2023mcdiff; wang2024motionctrl; shi2024motion; niu2024mofa; wu2024draganything; geng2024motion; li2024image; namekata2024sg; zhang2024tora), such as drag-based, trajectory-based, and optical flow-based techniques. Many of the existing motion control models yin2023dragnuwa; chen2023mcdiff; zhang2024tora require the complete pre-specified trajectory, specifying the location of the pixel on every generated frame. This reliance on full temporal information makes it difficult to use these models for simulation or prediction tasks.
@@
Motion Prompting (geng2024motion), a concurrent work, uses spatio-temporally sparse trajectories as a conditioning signal, enabling users to specify motion over a few frames for video extrapolation. While this might superficially resemble force control, crucial distinctions exist. First, global phenomena like wind or fluid dynamics are naturally expressed as forces but are difficult or impossible to represent with trajectories. Second, applied forces fundamentally depend on an object’s mass or material properties - a dependency absent when specifying motion or location (e.g., identical forces induce greater displacement in lighter objects). Third, specifying an object’s location across a few frames is not equivalent to an applied force; the same observed motion could result from numerous alternative causes, such as camera movement or internal object changes. We compare to Motion Prompting and demonstrate significantly better adherence to the conditioning force.
@@
Interactive world models: Paralleling the interest in video generation models, interactive world models (ha2018world) have gained significant attention. Despite extensive research in this area, investigations have predominantly concentrated on video game environments valevski2024diffusion; che2024gamegen; bruce2024genie. While a few contemporary studies have begun exploring real-world applications (e.g., bar2024navigation; agarwal2025cosmos), none explores interactions besides camera control or text. In contrast, our work focuses on interaction through physical forces.
@@
Physical simulators and hybrid approaches: Early work (davis2015visual; davis2015image) on generating video based on intuitive forces extracts modal bases of vibrating objects in 2D image space; these works, as well as their modern adaptations (li2024generative), represented motion as a series of vibrations with different frequencies and intensities, which works well for vibration-like motions but struggles to represent many types of motion, such as linear motion. This led to an alternative research direction explicitly incorporating physics solvers (chen2022virtual; zhong2024reconstruction; le2023differentiable; xie2024physgaussian; zhang2024physdreamer; huang2024dreamphysics; liu2024physics3d; lin2024phys4dgen; aira2024motioncraft). However, almost all of these techniques require the 3D geometry of the scenes. Recent work has focused on combining both physics simulators and generative models, trying to get the best of both worlds: accurate dynamics from the simulator and better appearance from generative models. For example, PhysGen (liu2024physgen) uses a rigid-body physics solver to model object collisions and then renders these scenes through a video generator, and PhysMotion (tan2024physmotion) uses a combination of a 3D physics solver and a video generation model. However, due to their usage of physics simulators, they are limited in the types of dynamics they can model. In contrast, we explore using the video generation as a simulator and do not use a physics simulator at inference time. We mention some concurrent works as well:  (li2025pisa) also explores the use of simulated videos to finetune generative models, but their focus is on modeling object freefall as opposed to learning physics-based control;  li2025wonderplay explores action-conditioned video generation, but their model requires the use of a physics simulator at inference time; and physctrl2025 explores force-conditioned video generation, but their framework requires learning a 3D point cloud trajectory model from synthetic data, and then passing that 4D temporal volume into a point cloud-conditioned video generation model.
@@ 3.1 Synthetic training data
To construct our global wind force dataset, we use a physics simulator to generate videos of flags waving in the wind. And we construct our local point force dataset in two parts: for the first part, we use a physics simulator to generate videos of a ball rolling across the ground; and for the second part, we use a model (zhang2024physdreamer) which integrates 3D Gaussians and a physics simulator to generate videos of a plant being poked. We provide more detail below.
Global force dataset: We use Blender to construct a dataset of flags waving in response to varying wind conditions. In order to generate a diverse dataset, we randomize multiple parameters for each video: flag quantity ($\mathrm{Unif}\{1,\dots,64\}$), flag color (from a set of 100), flag positions, camera placement, HDRIs (High Dynamic Range Images), which are 360-degree panoramic images used for lighting and background purposes (selected from 50 options on Polyhaven), wind direction in $[0,360)$, and wind speed in $[0,1]$, where 0 corresponds to no wind, and $1$ corresponds to very strong wind. Each video captures the flags’ transitions from stationary to wind-affected state. Our training dataset has 15k videos.
Local force dataset: The first scenario in our dataset comprises 12k videos of balls, with one of them rolling in response to being pushed by an unseen point-wise force (the force actor is not rendered), and the other balls remaining stationary. We generate these videos using Blender with randomized parameters: ball quantity ($\mathrm{Unif}\{2,3,4\}$), ball textures (soccer balls using a Polyhaven mesh [$p=2/3$] , or bowling balls modeled as smooth spheres [$p=1/3$]), ball colors (from a set of 108), ball positions, camera position, ground textures (from 42 Polyhaven options), target ball selection, force angle in $[0,360)$, and force magnitude in [0,1]. We assign the bowling ball to be four times the mass of the soccer ball with the goal of teaching the model mass-based dynamics. The second scenario (11k videos) utilizes PhysDreamer (zhang2024physdreamer), a generative-simulator hybrid, and features videos of a carnation swaying back and forth in response to being poked by an unseen force. We generate these videos with randomized camera position, contact points, force angles, and magnitudes. We use a mixed dataset with the goal of teaching the model that a point force can result in both simple linear motion, and complex oscillatory dynamics, depending on what type of object the force is applied to. In both scenarios, the force magnitude 0 corresponds to a very gentle poke, and the force magnitude $1$ corresponds to a much stronger poke.
@@ 3.2 Local and Global Force Prompts
Figure 3: Qualitative results for the Local Force (Poke) model. Top section: For local forces, the control signal can specify both the location, magnitude, and direction of the force. Bottom section: despite the limited training data, the model generalizes to different types of motion. We add blue lines to visualize a time-lapse of some objects’ movements.
@@
We note that force values across the ball rolling and plant poking training videos are not calibrated to any absolute physical scale. Instead, they follow intuitive relative physics where smaller force values (approaching $F=0$) correspond to gentle pokes resulting in minimal initial displacement, while larger force values produce stronger pokes with correspondingly greater initial displacement. We also wish to highlight that our force prompting models are fundamentally different from video generative models with trajectory-based control such as (zhang2024tora; geng2024motion). This is because the gaussian blob which serves as the force indicator for the point force model is generally far away from the pixels that it affects, as demonstrated in the complex oscillatory motion of the swaying flower in Figure 2. Similarly, the wind force control signal under-specifies which points must move to which locations, as that control signal is global and causal.
@@ 3.3 Architecture and Training
We build the force prompting models on top of CogVideoX-5B-I2V (yang2024cogvideox), a video generative model which accepts text and initial frame as conditional inputs. This model generates 49-frame videos at 8-fps. In order to integrate force prompt conditioning, we add a ControlNet (zhang2023adding) which inputs a physics control prompt $\pi$, processing it through downscaling, encoding, and temporal compression before combining with hidden states via a zero convolution. The ControlNet clones the first six transformer layers and fine-tunes them while keeping the base model’s transformer layers frozen. We base our implementation on (cogvideoxcontrolnetgithub) with modifications to adhere more closely to the original ControlNet design. We train the models on a four 80 GB A100 GPU cluster for 5000 training steps, which takes approximately one day. Training uses an instantaneous batch size per device of $1$, with two gradient accumulation steps, for an effective batch size of $8$. Full hyperparameter details are listed in Appendix A.1.
@@ 4 Quantitative and Qualitative Results
Baseline models: The first baseline is text-only, zero-shot, which uses the original CogVideoX model and describes the intended force with a string and appending it to the end of the original text prompt. Two example prompt string suffixes are “the apple is moved very forcefully, upwards and to the left”, and “the wind is medium strength, blowing right”. The second baseline is text-only, fine-tuned, which has the same ControlNet architecture as our force prompting model, but with zero-tensor control signals, as well as force suffixes added to the end of the text prompts during training. Our third baseline is Motion Prompting (geng2024motion), built on Lumiere (bar2024lumiere) (run by the authors). It is the only track-conditioned model that accepts temporally sparse tracks as conditioning signal. We simulate force prompt’s impulse by tracing push paths from target objects for the first 3 frames. While the model is meant to accept temporally sparse trajectories, 3 frames of trajectory is out of domain for the intended use case of Motion Prompting.
Human study for local force benchmark: We create a benchmark by curating 63 images from Pexels demonstrating three categories of physical interactions: 1) linear movement patterns (toy car, toy train on straight track, hot air balloon); 2) oscillatory movement patterns (windmill, pendulums, ornament, and swing); and 3) complex movement patterns (toy train on circular track, various plants including ivy, apple tree, and flowers). Table 1 presents human evaluation results for point forces, showing that despite training only on ball rolling (linear) and plant poking (complex) scenarios, our force prompting model demonstrates strong generalization across all motion categories. We note that this model successfully handles multiple forces “zero-shot” during inference, despite only being trained to handle a single force, as seen in Figure 3 and detailed in Appendix B.1.
@@
Human study comparing to PhysDreamer: The point force model, trained on data from a single carnation, demonstrates remarkable generalization to other plants, as we illustrate in Figures 1 and 3. To evaluate this generalization quantitatively, we compare our approach against PhysDreamer (zhang2024physdreamer), which employs 3D assets and an integrated physics simulator. Using their benchmark dataset of six plant species, our results in Table 2 show that the point force model successfully generalizes to various roses, tulips, and alocasia without specific training on these plants. While we do not claim to replace physics-based simulation approaches, our purely neural method offers exceptional generalizability and produces responses that align with “intuitive physics,” effectively conveying plausible physical interactions to human evaluators. We also also conduct an extended qualitative comparison with 8 other physics simulation models; see Appendix C.2 for more details.
@@ 5.1 Ablation Study #1: Composition of Synthetic Dataset
Global force training dataset ablation: For the global wind force task, we evaluate two diversity factors: flag quantity and background variety. We find that training with a single background leads to models that follow force physics but frequently fail to differentiate between foreground and background, reducing visual quality. Similarly, when restricting scenes to contain only one flag instead of a variable number ($\mathrm{Unif}\{1,\dots,64\}$), the model successfully models cloth mechanics but fails to generalize to other materials. In these cases, smoke from campfires remains unaffected by wind, and confetti either doesn’t respond or stays unnaturally suspended. We also observe that bubbles don’t respond to wind, while human limbs incorrectly billow like cloth. These failures indicate that insufficient scene diversity causes the model to overfit to stationary backgrounds and limited material interactions. These findings are illustrated in Figure 7 as well as the project webpage. Additionally, we trained a unified model to learn both point force prompts and wind force prompts. We found that this results in more dynamic backgrounds, but has slightly less robust point force control. Additional details are in Appendix C.3.